# Contrastive Preference Optimization: Pushing the Boundaries of LLM Performance in Machine Translation

### 18 Jan 2024 v2

@@ Abstract
Moderate-sized large language models (LLMs) – those with 7B or 13B parameters – exhibit promising machine translation (MT) performance. However, even the top-performing 13B LLM-based translation models, like ALMA (Xu et al., 2023), does not match the performance of state-of-the-art conventional encoder-decoder translation models or larger-scale LLMs such as GPT-4 (OpenAI, 2023). In this study, we bridge this performance gap. We first assess the shortcomings of supervised fine-tuning for LLMs in the MT task, emphasizing the quality issues present in the reference data, despite being human-generated. Then, in contrast to supervised fine-tuning which mimics reference translations, we introduce Contrastive Preference Optimization (CPO), a novel approach that trains models to avoid generating adequate but not perfect translations. Applying CPO to ALMA models with only 22K parallel sentences and 0.1% parameters yields significant improvements. The resulting model, called ALMA-R, can match or exceed the performance of the WMT competition winners and GPT-4 on WMT’21, WMT’22 and WMT’23 test datasets.111We release our code and models at: https://github.com/fe1ixxu/ALMA.
@@ 1 Introduction
Machine translation (MT) predominantly utilizes transformer encoder-decoder architectures (Vaswani et al., 2017), which is evident in prominent models such as NLLB-200 (NLLB TEAM et al., 2022), M2M100 (Fan et al., 2021), BIBERT (Xu et al., 2021), and MT5 (Xue et al., 2021). However, the emergence of decoder-only large language models (LLMs) such as the GPT series (Brown et al., 2020; OpenAI, 2023), Mistral (Jiang et al., 2023), LLaMA (Touvron et al., 2023a, b), Falcon (Almazrouei et al., 2023), and others, which have shown remarkable efficacy in various NLP tasks, which attracts the interest of developing machine translation with these decoder-only LLMs. Recent studies (Zhu et al., 2023a; Jiao et al., 2023b; Hendy et al., 2023; Kocmi et al., 2023; Freitag et al., 2023) indicate that larger LLMs such as GPT-3.5 (175B) and GPT-4 exhibit strong translation abilities. However, the performance of smaller-sized LLMs (7B or 13B) still falls short when compared to conventional translation models (Zhu et al., 2023a).
Therefore, there are studies intend to enhance the translation performance for these smaller LLMs (Yang et al., 2023; Zeng et al., 2023; Chen et al., 2023; Zhu et al., 2023b; Li et al., 2023; Jiao et al., 2023a; Zhang et al., 2023), but their improvements are relatively modest, primarily due to the predominant pre-training of LLMs on English-centric datasets, resulting in limited linguistic diversity (Xu et al., 2023). Addressing this limitation, Xu et al. (2023) introduce ALMA models, which initially fine-tune LLaMA-2 (Touvron et al., 2023b) with extensive monolingual data in various languages to enhance their multilingual abilities, and then perform supervised fine-tune (SFT) with small but high-quality parallel data to induce the model toward the translation generation. This method has allowed ALMA models to outperform all prior moderated-size LLMs, and even larger models such as GPT-3.5, in the translation task. Nonetheless, the performance still lags behind leading translation models such as GPT-4 and WMT competition winners. Our study bridges this gap by further fine-tuning ALMA models with our novel training method Contrastive Preference Optimization (CPO) and minimal costs, i.e., only 12M learnable parameters (equivalent to 0.1% of the original model size) and a 22K dataset for 10 directions. The fine-tuned model is referred to as ALMA-R. A detailed performance comparison is illustrated in Figure 1.
@@ 3 Contrastive Preference Optimization
In this section, we present a novel preference learning technique, termed Contrastive Preference Optimization (CPO). This method is designed to guide the model in developing a propensity for generating ‘better’ translations while simultaneously learning to reject ’worse’ ones, even in cases where these ‘worse’ translations are of high quality but not perfect.
@@ 3.1 Deriving the CPO Objective
where is a pre-trained language (translation) model , is the Sigmoid function, and is a hyperparameter. The DPO loss is derived a reparameterization process of the ground-truth reward and the corresponding optimal policy in the Proximal Policy Optimization (PPO) framework (Schulman et al., 2017). As a result, DPO training can be conducted in a supervised fine-tuning style, as it relies exclusively on labeled preference data and does not require interaction between agents and their environment.
@@
Initially, we set , representing an ideal policy that perfectly aligns the true data distribution of the preferred data. Specifically, for any given data point from the dataset , the conditions and hold true. This contrasts with the conventional approach of assigning to the initial pre-trained model checkpoint. This modification is feasible because the primary role of is to prevent deviation from the original model, and our approach similarly aligns with this goal, maintaining the model’s proximity to the ideal preferred model. Consequently, under this setup, the predictions for preferred data do not require reweighting by the reference model, and Equation 2 can be reformulated as follows :
However, is unknown as the ideal model is unreachable, but we can approximate the optimization. After expanding the Sigmoid function and removing the non-parameterized term, the loss becomes to:
Considering that , the loss can be upper bounded as follows:
Upon restructuring the above loss into a Sigmoid function, we derive our new preference learning objective without computing by minimizing the upper bound of Equation 3.1:
A detailed step-by-step derivation from Equation 3 to 6 is provided in Appendix C. Our approach also involves implementing a straightforward signal to guide the learnable policy towards the preferred data distribution. Specifically, we incorporate a log-likelihood supervised fine-tuning loss applied to the preferred data:
By combing Equation 6 and 7, our CPO loss is finally formulated as follows:
The transition from DPO to CPO resolves issues of memory and speed inefficiency. This efficiency is achieved as CPO only requires the storage and processing of one policy , so CPO facilitates preference learning with the same space and time complexity as common SFT methods. We also show the substantially superior performance of CPO compared to DPO in Section 4.
3.2 Triplet Preference Data
We here describe the construction of our preference data . This dataset is developed using the FLORES-200 data (both development and test sets) and encompasses the same language pairs as discussed in Section 2. For each language pair, the dataset comprises 2009 parallel sentences.
@@ 4.4 WMT’21 and WMT’22 Results
Comparing With SFT and DPO All training objectives in our study are fine-tuned using the ALMA-13B-LoRA model as a base. In Table 2 and 3, we employ deep green boxes to indicate improvements exceeding 1 point, and shallow green boxes for improvements less than 1 point. Conversely, red boxes denote a decline in performance. We observe that SFT on preferred data marginally enhances the ALMA model’s translation capability for xxen, and results in a slight deterioration for enxx. Similarly, DPO slightly decreases model performance. In contrast, CPO demonstrates significant improvements across all translation directions.
@@ 4.5 WMT’23 Results
We show the average results across all six directions in Table 4, and provide the performance in each direction in Appendix F due to the space constraints. Consistent with observations from WMT’21 and WMT’22, ALMA-13B-R surpasses contemporary moderate-size LLM-based translators such as ALMA-13B-LoRA and TowerInstruct, and either matches or exceeds the performance of WMT winners.
| Avg. | |||
@@
5 Analyses
All analyses were conducted using the WMT’21 and WMT’22 test datasets, with their averaged performance being reported.
5.1 Are Translations Really Better or Just Metric-Preferred?
In our study, since the preferred data is selected by reference-free models and the same models are used for evaluation, we investigate the potential for ’cheating’ in the scoring process. Specifically, we question whether the improved translation scores reflect genuinely better translations or if they simply align more closely with the evaluation model’s preferences? This inquiry is addressed in two parts:
At the metric level, we examine if training a model on data preferred by a specific metric (such as KIWI-XXL) yields improvements that are consistent across other metrics. To investigate this, we reconstruct the preference data using only KIWI-XXL or XCOMET and re-train the ALMA-13B-LoRA model using the CPO method. The results, presented in Table 5, do not indicate a significant bias towards the metric used for selecting preferred data. We observed similar and consistent improvements across all metrics, regardless of the specific metric used to select the preferred data.
| Loss Objective | KIWI-22 | KIWI-XXL | XCOMET |
| Translating to English (xxen) | |||
| 80.51 | 81.35 | 86.58 | |
| 80.96 | 81.99 | 88.40 | |
| 81.33 | 82.43 | 89.11 | |
| Translating from English (enxx) | |||
| 82.23 | 82.06 | 92.22 | |
| 82.57 | 82.42 | 92.54 | |
| 83.34 | 85.74 | 94.05 | |
@@ 5.2 Ablation Study
CPO Loss Components The CPO loss function consists of two components: for preference learning, and , which ensures the model does not deviate significantly from the preferred data distribution. To illustrate the significance of each term, we re-train the model exclusively with one of the components. It is important to note that training solely with equates to the baseline scenario of SFT on preferred data. As depicted in Table 6, the inclusion of both terms yields the optimal performance, while the absence of either leads to a decrease in performance.
Preference Data Components: Our preference data selection involves choosing preferred and dis-preferred translations from a triplet consisting of outputs from GPT-4, ALMA, and the gold reference. In Table 7, we emphasize the significance of the data generated by both ALMA and GPT-4. The results indicate a notable decline in performance when ALMA data is excluded, particularly in the enxx direction, with the KIWI-XXL score dropping from 85.74 to 84.50 on average. Conversely, omitting GPT-4 data leads to a significant performance decrease in the xxen direction. This demonstrates that data generated by both systems plays a helpful role in enhancing model performance.
| Preference Data | KIWI-22 | KIWI-XXL | XCOMET |
| Translating to English (xxen) | |||
| ALMA + Ref | 81.10 | 82.21 | 87.80 |
| GPT-4 + Ref | 81.20 | 82.28 | 89.12 |
| All of them | 81.33 | 82.43 | 89.11 |
| Translating from English (enxx) | |||
| ALMA + Ref | 82.58 | 85.30 | 93.09 |
| GPT-4 + Ref | 82.37 | 84.50 | 93.10 |
| All of them | 83.34 | 85.74 | 94.05 |
@@
5.3 Does The Quality of Dis-preferred Data Matter?
In this new dataset, the preferred translation () remains the best of the three translation candidates, selected in the same manner as in Section 3.2. However, the dis-preferred translation is intentionally modified to be a noised version of . We applied random deletions of words with a probability of 0.15 and word swaps within a range of 1 with a probability of 0.3, following the method suggested by Zeng et al. (2023) for creating manually noised dis-preferred data. This approach produces worse translations that are unnatural and artificial.
@@ 6 Conclusion
In this study, we initially proposed the potential quality issues of gold references in machine translation tasks, highlighting instances where advanced translation models outperform these references. This finding challenges the conventional assumption of gold references as the optimal standard, impacting not only model training, which often relies on minimizing the difference between predicted tokens and gold references, but also potentially skewing results in reference-based evaluation metrics. Subsequently, we introduce Contrastive Preference Optimization, a more efficient variant of of DPO. This method leverages both model-generated and reference data to guide the model in avoiding near-perfect yet flawed translations and learning superior ones. Our developed model, ALMA-13B-R, stands out as the first moderate-size LLM-based translation model to match, and in some cases surpass, the performance of cutting-edge systems such as GPT-4 and WMT competition winners, marking a significant advancement in the field of neural machine translation.
@@ Acknowledgements
We extend our gratitude to Hieu Hoang, Marcin Junczys-Dowmunt, Huda Khayrallah, Thamme Gowda, Vikas Raunak, Matt Post, Anoop Kunchukuttan, Roman Grundkiewicz, Philipp Koehn, Hany Hassan Awadalla, Arul Menezes and Vishal Chowdhary for their insightful and valuable suggestions.
@@ Appendix A Comprehensive Results of WMT’21 And WMT’22
We show the comprehensive results for enxx In Table 9 and xxen in Table 10. In this section, our study additionally includes results from recently released LLM-based translators, including Bayling-13B (Zhang et al., 2023), BigTranslate (Yang et al., 2023), ALMA-13B-LoRA (Xu et al., 2023), the zero-shot performances of LLaMA-1-13B (Touvron et al., 2023a) and LLaMA-2-13B (Touvron et al., 2023b). We also compare these with the most advanced current translation models, such as WMT competition winners, GPT-4, GPT-3.5-text-davinci-003, Google Translate, and NLLB-3.3B. Importantly, we also present the performance of ALMA-7B-R here, which is fine-tuning on AMLA-7B-LoRA with CPO method. Except for reference-free evaluation, we also report two commonly used reference-based metrics, sacreBLEU (Post, 2018; Papineni et al., 2002) and COMET-22 (Unbabel/wmt22-comet-da) (Rei et al., 2022), in our analysis.
@@
Introducing ALMA-7B-R In this study, we extend the ALMA-13B-R training methodology to a 7B model size, specifically fine-tuning ALMA-7B-LoRA using the CPO method with the same preference data as ALMA-13B-R. Consistent with our findings from ALMA-13B-R, the application of CPO significantly enhances performance. The ALMA-7B-R model surpasses ALMA-7B-LoRA by an average of 2.54 KIWI-XXL points and 1.44 XCOMET points for enxx translations, and by 0.89 KIWI-XXL points and 3.69 XCOMET points for xxen.
@@
Comparing with Advanced Translation Models Our model, ALMA-13B-R, is benchmarked against the most advanced current models, demonstrating performance comparable to GPT-4 and WMT winners. It surpasses leading commercial translation tools such as Google Translate in many cases and top multilingual translation models like NLLB and GPT-3.5. Stop Using BLEU BLEU, a metric extensively utilized for decades, often diverges from neural-based and reference-free metrics, a phenomenon also observed in previous studies (Xu et al., 2023; Freitag et al., 2023). For instance, WMT competition winners often exhibit superior performance according to BLEU (or COMET-22), yet this is not corroborated by reference-free models. A case in point is the WMT winners scoring an exceptionally high 64.14 BLEU in csen translations, significantly outperforming other models by 20 BLEU points. However, reference-free evaluations suggest these translations are inferior to those generated by our models and GPT-4. We hypothesize that this discrepancy may arise from WMT models being trained on domain-specific data closely related to the WMT test set, leading to high lexical matches but lacking semantic depth as evaluated by neural-based metrics. While BLEU scores are effective for assessing basic functionality in weaker models, their utility diminishes with advanced translation models capable of generating diverse translations. In such contexts, relying solely on BLEU for evaluation appears increasingly outdated.
@@
Towards Reference-Free Metrics Neural-based, reference-dependent metrics like COMET-22 demonstrate greater consistency with reference-free metrics and robustness compared to BLEU. For instance, with COMET-22, our models show significant improvements like other reference-free models over ALMA-13B-LoRA and comparable performance to GPT-4, e.g., 87.74 (Ours) vs. 87.68 (GPT-4) when enxx. However, it is important to note that, according to reference-free metrics, gold references are often inferior to system-generated translations, potentially indicating quality issues in the references that could impact COMET-22 evaluations. Consequently, inconsistencies still exist between COMET-22 and reference-free models like XCOMET. For example, XCOMET rates our ALMA+CPO model on average higher than WMT winners (89.11 vs. 87.13), while COMET-22 favors WMT winners (85.21 vs. 85.60). In line with the recommendations in Freitag et al. (2023), we advocate for the use of reference-free models to circumvent the potential quality issues of references.

### 02 Feb 2024 v3

@@ Abstract
Moderate-sized large language models (LLMs) – those with 7B or 13B parameters – exhibit promising machine translation (MT) performance. However, they do not match the performance of state-of-the-art conventional encoder-decoder translation models or larger-scale LLMs such as GPT-4 (OpenAI, 2023). In this study, we bridge this performance gap. We first assess the shortcomings of supervised fine-tuning for LLMs in the MT task, emphasizing the quality issues present in the reference data, despite being human-generated. Then, in contrast to supervised fine-tuning which mimics reference translations, we introduce Contrastive Preference Optimization (CPO), a novel approach that trains models to avoid generating adequate but not perfect translations. Applying CPO to ALMA (Xu et al., 2023) models with only 22K parallel sentences and 0.1% parameters yields significant improvements. The resulting model, called ALMA-R, can match or exceed the performance of the WMT competition winners and GPT-4 on WMT’21, WMT’22 and WMT’23 test datasets.111We release our code and models at: https://github.com/fe1ixxu/ALMA.
@@ 1 Introduction
Machine translation (MT) predominantly utilizes transformer encoder-decoder architectures (Vaswani et al., 2017), which is evident in prominent models such as NLLB-200 (NLLB TEAM et al., 2022), M2M100 (Fan et al., 2021), BIBERT (Xu et al., 2021), and MT5 (Xue et al., 2021). However, the emergence of decoder-only large language models (LLMs) such as the GPT series (Brown et al., 2020; OpenAI, 2023), Mistral (Jiang et al., 2023), LLaMA (Touvron et al., 2023a, b), Falcon (Almazrouei et al., 2023), inter alia, which have shown remarkable efficacy in various NLP tasks, which attracts the interest of developing machine translation with these decoder-only LLMs. Recent studies (Zhu et al., 2023a; Jiao et al., 2023b; Hendy et al., 2023; Kocmi et al., 2023; Freitag et al., 2023) indicate that larger LLMs such as GPT-3.5 (175B) and GPT-4 exhibit strong translation abilities. However, the performance of smaller-sized LLMs (7B or 13B) still falls short when compared to conventional translation models (Zhu et al., 2023a).
Therefore, there are studies intend to enhance the translation performance for these smaller LLMs (Yang et al., 2023; Zeng et al., 2023; Chen et al., 2023; Zhu et al., 2023b; Li et al., 2023; Jiao et al., 2023a; Zhang et al., 2023), but their improvements are relatively modest, primarily due to the predominant pre-training of LLMs on English-centric datasets, resulting in limited linguistic diversity (Xu et al., 2023). Addressing this limitation, Xu et al. (2023) initially fine-tune LLaMA-2 (Touvron et al., 2023b) with extensive non-English monolingual data to enhance their multilingual abilities, and then perform supervised fine-tune (SFT) with high-quality parallel data to instruct the model to generate translations. Their model, named ALMA, outperforms all prior moderated-size LLMs, and even larger models such as GPT-3.5, in the translation task. Nonetheless, the performance still lags behind leading translation models such as GPT-4 and WMT competition winners. Our study bridges this gap by further fine-tuning ALMA models with our novel training method Contrastive Preference Optimization (CPO) and minimal costs, i.e., only 12M learnable parameters (equivalent to 0.1% of the original model size) and a 22K dataset for 10 directions. The fine-tuned model is referred to as ALMA-R. A detailed performance comparison is illustrated in Figure 1.
@@
Preference Data We build and release a high-quality preference data for the machine translation area.
@@ 3 Contrastive Preference Optimization
To learn an objective that fosters superior translations and rejects inferior ones, access to labeled preference data is essential, yet such data is scarce in machine translation. In this section, we first describe the construction of our preference data and then introduces a preference learning technique, contrastive preference optimization (CPO).
3.1 Triplet Preference Data
We here details our methodology for constructing preference data . This dataset is developed using the FLORES-200 data (both development and test sets) and encompasses the same language pairs as discussed in Section 2. For each language pair, the dataset comprises 2009 parallel sentences.
@@ 3.2 Deriving the CPO Objective
where is a pre-trained language (translation) model , is the Sigmoid function, and is a hyperparameter. DPO training can be conducted in a supervised fine-tuning style, as it relies exclusively on labeled preference data and does not require interaction between agents and their environment.
@@
The memory- or speed- inefficiency can be resolved when is set as a uniform prior , as the terms and cancel each other out. This negates the need for additional computations and storage beyond the policy model itself. Thus, we initially demonstrate that the DPO loss can be effectively approximated using a uniform reference model:
Specifically, we prove the below Theorem in Appendix C.
Theorem 1.
When is defined as , an ideal policy that precisely aligns with the true data distribution of preferred data, the DPO loss is upper bounded by , where is a constant.
The approximation in Equation 3 is effective because it minimizes the upper boundary of the DPO loss. The proof relies on an important assumption of . Contrary to common practices where is set as the initial SFT checkpoint, our approach considers it as the ideal policy we aim to reach. Although the ideal policy is unknown and unattainable during model training, it is not engaged in the loss after our approximation.
Furthermore, we incorporate a behavior cloning (BC) regularizer (Hejna et al., 2023) to ensure that does not deviate from the preferred data distribution:
where is a small positive constant and is Kullback–Leibler (KL) divergence. The regularizer can boil down to adding a SFT term on the preferred data (a detailed explanation is provided in Appendix C):
The above is the formulation of our CPO loss, which includes one preference learning term and one negative log likelihood term .
@@ 4.4 WMT’21 and WMT’22 Results
Comparing With SFT and DPO All training objectives in our study are fine-tuned using the ALMA-13B-LoRA model as a base. In Table 2 and 3, we observe that SFT on preferred data marginally enhances the ALMA model’s translation capability for xxen, and results in a slight deterioration for enxx. Similarly, DPO slightly decreases model performance. In contrast, CPO demonstrates significant improvements across all translation directions.
@@ 4.5 WMT’23 Results
We show the average results across all six directions in Table 4, and provide the performance in each direction in Appendix G due to the space constraint. Consistent with observations from WMT’21 and WMT’22, ALMA-13B-R surpasses contemporary moderate-size LLM-based translators such as ALMA-13B-LoRA and TowerInstruct, and either matches or exceeds WMT winners.
@@
5 Analyses
All analyses use the WMT’21 and WMT’22 test sets, with their averaged performance being reported.
5.1 Are Translations Really Better or Just Metric-Preferred?
In our study, since the preferred data is selected by reference-free models and the same models are used for evaluation, we investigate the potential for ’cheating’ in the scoring process. Specifically, we question whether the improved translation scores reflect genuinely better translations or if they simply align more closely with the evaluation model’s preferences. This inquiry is addressed in two parts:
At the metric level, we examine if training a model on data preferred by a specific metric (such as KIWI-XXL) yields improvements that are consistent across other metrics. To investigate this, we reconstruct the preference data using only KIWI-XXL or XCOMET and re-train the ALMA-13B-LoRA model using the CPO method. The results, presented in Table 5, do not indicate a significant bias towards the metric used for selecting preferred data. We observed similar and consistent improvements across all metrics, regardless of the specific metric used to select the preferred data. Considering Comet-series models may be positive correlated, we further evaluate ALMA-R using a non-comet metric, BLEURT (Sellam et al., 2020), and also observe significant improvements in Appendix H. The inclusion of a third-party evaluation metric further substantiates the superior translation quality of ALMA-R.
@@ 5.2 Ablation Study
CPO Loss Components The CPO loss function consists of two components: for preference learning, and , which ensures the model does not deviate significantly from the preferred data distribution. To illustrate the significance of each term, we re-train the model exclusively with one of the components. It is important to note that training solely with equates to the baseline scenario of SFT on preferred data. As depicted in the left of Figure 4, the inclusion of both terms yields the optimal performance, while the absence of either leads to a decrease in performance. In Appendix I, we also show that incorporating into the DPO loss yields significant improvements.
Preference Data Components: Our preference data selection involves choosing preferred and dis-preferred translations from a triplet consisting of outputs from GPT-4, ALMA, and the gold reference. In the right of Figure 4, we emphasize the significance of the data generated by both ALMA and GPT-4. The results indicate a notable decline in performance when ALMA data is excluded in the enxx direction. Conversely, omitting GPT-4 data leads to a significant performance decrease in the xxen direction. This demonstrates that data generated by both systems plays a helpful role in enhancing model performance.
5.3 Does The Quality of Dis-preferred Data Matter?
In this new dataset, the preferred translation () remains the best of the three translation candidates, selected in the same manner as in Section 3.1. However, the dis-preferred translation is intentionally modified to be a noised version of . We applied random deletions of words with a probability of 0.15 and word swaps within a range of 1 with a probability of 0.3, following the method suggested by Zeng et al. (2023) for creating manually noised dis-preferred data. This approach produces worse translations that are artificial.
@@ 6 Conclusion
In this study, we initially proposed the potential quality issues of gold references in the MT task, highlighting instances where advanced translation models can outperform these references. This finding not only challenges model training via SFT, but also the evaluation procedure that uses reference-based metrics. Subsequently, we introduce Contrastive Preference Optimization, a more efficient variant of of DPO. This method leverages both model-generated and reference data to guide the model in avoiding near-perfect yet flawed translations and learning superior ones. Our developed model, ALMA-13B-R, stands out as the first moderate-size LLM-based translation model to match, and in some cases surpass, the performance of GPT-4 and WMT competition winners, marking a significant advancement in the field of MT.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Translation. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgements
We express our profound appreciation to Hieu Hoang, Marcin Junczys-Dowmunt, Huda Khayrallah, Thamme Gowda, Vikas Raunak, Matt Post, Anoop Kunchukuttan, Roman Grundkiewicz, Philipp Koehn, Hany Hassan Awadalla, Arul Menezes, and Vishal Chowdhary for their engaging and valuable discussions that greatly enriched our work. Special thanks to Tom Kocmi for his innovative suggestion to enhance numerical data visibility using a dynamic threshold determined by estimated accuracy. Our gratitude also extends to Pushpendre Rastogi and Joey Hejna for their insightful recommendations on the CPO theory. Furthermore, we acknowledge the Unbabel Team for their valuable advice on incorporating non-COMET metrics into our analysis.
Appendix Contents
| Appendix Sections | Contents | |
| Appendix A |
| Appendix B |
| Appendix C |
| Appendix D |
| Appendix E |
| Appendix F |
| Appendix G |
| Appendix H |
| Appendix I |
@@ Appendix A Comprehensive Results of WMT’21 And WMT’22
We show the comprehensive results for enxx In Table 7 and xxen in Table 8. In this section, our study additionally includes results from recently released LLM-based translators, including Bayling-13B (Zhang et al., 2023), BigTranslate (Yang et al., 2023), ALMA-13B-LoRA (Xu et al., 2023), the zero-shot performances of LLaMA-1-13B (Touvron et al., 2023a) and LLaMA-2-13B (Touvron et al., 2023b). We also compare these with the most advanced current translation models, such as WMT competition winners, GPT-4, GPT-3.5-text-davinci-003, Google Translate, NLLB-3.3B, and MADLAD-10B (Kudugunta et al., 2023). Importantly, we also present the performance of ALMA-7B-R here, which is fine-tuning on AMLA-7B-LoRA with CPO method. Except for reference-free evaluation, we also report two commonly used reference-based metrics, sacreBLEU (Post, 2018; Papineni et al., 2002) and COMET-22 (Rei et al., 2022).
@@
Introducing ALMA-7B-R In this study, we extend the ALMA-13B-R training methodology to a 7B model size, specifically fine-tuning ALMA-7B-LoRA using the CPO method with the same preference data as ALMA-13B-R. Consistent with our findings from ALMA-13B-R, the application of CPO significantly enhances performance.
Comparing with Advanced Translation Models Our model, ALMA-13B-R, is benchmarked against the most advanced current models, demonstrating performance comparable to GPT-4 and WMT winners. It surpasses leading commercial translation tools such as Google Translate in many cases and top multilingual translation models like NLLB, MADLAD-10B and GPT-3.5.
@@
Stop Using BLEU BLEU, a metric extensively utilized for decades, often diverges from neural-based and reference-free metrics, a phenomenon also observed in previous studies (Xu et al., 2023; Freitag et al., 2023). For instance, WMT competition winners often exhibit superior performance according to BLEU (or COMET-22), yet this is not corroborated by reference-free models. A case in point is the WMT winners scoring an exceptionally high 64.14 BLEU in csen translations, significantly outperforming other models by 20 BLEU points. However, reference-free evaluations suggest these translations are inferior to those generated by our models and GPT-4. We hypothesize that this discrepancy may arise from WMT models being trained on domain-specific data closely related to the WMT test set, leading to high lexical matches but lacking semantic depth as evaluated by neural-based metrics. While BLEU scores are effective for assessing basic functionality in weaker models, their utility diminishes with advanced translation models capable of generating diverse translations. In such contexts, relying solely on BLEU for evaluation appears increasingly outdated.
@@
Towards Reference-Free Metrics Neural-based, reference-dependent metrics like COMET-22 demonstrate greater consistency with reference-free metrics and robustness compared to BLEU. For instance, with COMET-22, our models show significant improvements like other reference-free models over ALMA-13B-LoRA and comparable performance to GPT-4, e.g., 87.74 (Ours) vs. 87.68 (GPT-4) when enxx. However, it is important to note that, according to reference-free metrics, gold references are often inferior to system-generated translations, potentially indicating quality issues in the references that could impact COMET-22 evaluations. Consequently, inconsistencies still exist between COMET-22 and reference-free models like XCOMET. For example, XCOMET rates ALMA-R model on average higher than WMT winners (89.11 vs. 87.13), while COMET-22 favors WMT winners (85.21 vs. 85.60). In line with the recommendations in Freitag et al. (2023), we advocate for the use of reference-free models to circumvent the potential quality issues of references.